Rotational motion with a constant nonzero acceleration is not uncommon in the world around us. For instance, many machines have spinning parts. When the machine is turned on or off, the spinning parts tend to change the rate of their rotation with virtually constant angular acceleration. Many introductory problems in rotational kinematics involve motion of a particle with constant, nonzero angular acceleration. The kinematic equations for such motion can be written as

\theta (t) = \theta_0 +\omega_0t + \frac{1}{2}\alpha t^2

and

\omega (t) = \omega_0 + \alpha t.

Here, the symbols are defined as follows:

  • theta(t) is the angular position of the particle at time t.
  • theta_0 is the initial angular position of the particle.
  • omega(t) is the angular velocity of the particle at time t.
  • omega_0 is the initial angular velocity of the particle.
  • alpha is the angular acceleration of the particle.
  • t is the time that has elapsed since the particle was located at its initial position.
In answering the following questions, assume that the angular acceleration is constant and nonzero: \alpha \neq 0.
A)
True or false: The quantity represented by theta is a function of time (i.e., is not constant).
true
false
B)
True or false: The quantity represented by theta_0 is a function of time (i.e., is not constant).
true
false
C)
True or false: The quantity represented by omega_0 is a function of time (i.e., is not constant).
true
false
D)
True or false: The quantity represented by omega is a function of time (i.e., is not constant).
true
false
E)
Which of the following equations is not an explicit function of time t? Keep in mind that an equation that is an explicit function of time involves t as a variable.
\theta=\theta_0+\omega_0 t+\frac{1}{2}\alpha t^2
\omega=\omega_0+\alpha t
\omega^2=\omega_0^2+2\alpha(\theta-\theta_0)
F)
In the equation \omega=\omega_0+\alpha t, what does the time variable t represent?
Choose the answer that is always true. Several of the statements may be true in a particular problem, but only one is always true.
the moment in time at which the angular velocity equals \omega_0
the moment in time at which the angular velocity equals \omega
the time elapsed from when the angular velocity equals \omega_0 until the angular velocity equals \omega
G)
Consider two particles A and B. The angular position of particle A, with constant angular acceleration, depends on time according to \theta_{\rm A}(t)=\theta_0+\omega_0t+\frac{_1}{^2}\alpha t^2. At time t=t_1, particle B, which also undergoes constant angular acceleration, has twice the angular acceleration, half the angular velocity, and the same angular position that particle A had at time t=0.
Which of the following equations describes the angular position of particle B?
\theta_{\rm B}(t)=\theta_0+2\omega_0t+\frac{1}{4}\alpha t^2
\theta_{\rm B}(t)=\theta_0+\frac{1}{2}\omega_0t+\alpha t^2
\theta_{\rm B}(t)=\theta_0+2\omega_0(t-t_1) +\frac{1}{4}\alpha (t-t_1)^2
\theta_{\rm B}(t)=\theta_0+\frac{1}{2}\omega_0(t-t_1)+\alpha (t-t_1)^2
\theta_{\rm B}(t)=\theta_0+2\omega_0(t+t_1) +\frac{1}{4}\alpha (t+t_1)^2
\theta_{\rm B}(t)=\theta_0+\frac{1}{2}\omega_0(t+t_1)+\alpha (t+t_1)^2
H)
How long after the time t_1 does the angular velocity of particle B equal that of particle A?
\frac{\omega_0}{4\alpha}
\frac{\omega_0+4\alpha t_1}{2\alpha}
\frac{\omega_0+2\alpha t_1}{2\alpha}
The two particles never have the same angular velocity.
Answer #1
Concepts and reason

The concept required to solve the given problem is rotational equations of motion.

Use the rotational equations of motion together with the definition of angular velocity and displacement to derive the various conclusions.

Fundamentals

The equations of motion for rotational motion are,

ω2=ω02+2αθ{\omega ^2} = {\omega _0}^2 + 2\alpha \theta

ω=ω0+αt\omega = {\omega _0} + \alpha t

θ=ω0t+12αt2\theta = {\omega _0}t + \frac{1}{2}\alpha {t^2}

Here, ω\omega is the final velocity, ω0{\omega _0} is the initial velocity, α\alpha is the angular acceleration, tt is the time and θ\theta is the angular displacement.

(A)

The formula for θ\theta is given to be,

θ(t)=θ0+ω0t+12αt2\theta \left( t \right) = {\theta _0} + {\omega _0}t + \frac{1}{2}\alpha {t^2}

Here, θ0{\theta _0} is the initial displacement, tt is the time, ω0{\omega _0} is the initial velocity and α\alpha is the angular acceleration.

Since it is clearly seen that the formula for angular displacement at time tt contains tandt2t{\rm{ and }}{t^2} terms, the quantity represented by θ\theta is a function of time.

(B)

The quantity θ0{\theta _0} is the initial angular position of the particle which is a constant and hence will not change with time.

Thus, the quantity θ0{\theta _0} is independent on time.

(C)

The quantity ω0{\omega _0} is the initial angular velocity of the particle which is a constant and hence will not change with time.

Thus, the quantity ω0{\omega _0} is independent on time.

(D)

The formula for ω\omega is given to be,

ω(t)=ω0+αt\omega \left( t \right) = {\omega _0} + \alpha t

Here, tt is the time, ω0{\omega _0} is the initial velocity and α\alpha is the angular acceleration.

Since it is clearly seen that the formula for angular velocity at time tt contains the term tt. Thus, quantity represented by ω\omega is a function of time.

(E)

The formula for θ\theta is given as,

θ(t)=θ0+ω0t+12αt2\theta \left( t \right) = {\theta _0} + {\omega _0}t + \frac{1}{2}\alpha {t^2}

Since it is clearly seen that the formula for angular displacement at time tt contains tandt2t{\rm{ and }}{t^2} terms, thus, the equation θ(t)=θ0+ω0t+12αt2\theta \left( t \right) = {\theta _0} + {\omega _0}t + \frac{1}{2}\alpha {t^2}is a function of time.

The formula for ω\omega is given to be,

ω(t)=ω0+αt\omega \left( t \right) = {\omega _0} + \alpha t

Since it is clearly seen that the formula for angular velocity at time tt contains the term tt, thus, the equation ω(t)=ω0+αt\omega \left( t \right) = {\omega _0} + \alpha t is a function of time.

The formula for ω\omega is also given to be,

ω2=ω02+2α(θθ0){\omega ^2} = {\omega _0}^2 + 2\alpha \left( {\theta - {\theta _0}} \right)

Since it is clearly seen that the formula for angular velocity at time tt does not contain the term tt, hence, the equation ω2=ω02+2α(θθ0){\omega ^2} = {\omega _0}^2 + 2\alpha \left( {\theta - {\theta _0}} \right)is not an explicit function of time.

(F)

The formula for ω\omega is given to be,

ω(t)=ω0+αt\omega \left( t \right) = {\omega _0} + \alpha t

Here, the variable tt represents the time elapsed from when the angular velocity equalsω0{\omega _0} until the angular velocity equals ω\omega . It is the time taken by the particle to reach its final angular velocity ω\omega starting from initial angular velocityω0{\omega _0}.

(G)

The angular displacement of the particle A at t=0st = 0{\rm{ s}} is,

θA(t)=θ0+ω0t+12αt2{\theta _{\rm{A}}}\left( t \right) = {\theta _0} + {\omega _0}t + \frac{1}{2}\alpha {t^2}

At time t=t1t = {t_1}, angular velocity of particle B is,

ωB=12ω0{\omega _B} = \frac{1}{2}{\omega _0}

Angular acceleration of particle B is,

αB=2α{\alpha _B} = 2\alpha

The angular displacement of the particle B at tt will be,

θB(t)=θ0+ω0t+12αBt2{\theta _{\rm{B}}}\left( t \right) = {\theta _0} + {\omega _0}t' + \frac{1}{2}{\alpha _{\rm{B}}}{t'^2}

Substitute tt1t - {t_1} for tt', 12ω0\frac{1}{2}{\omega _0} for ωB{\omega _B}and 2α2\alpha for αB{\alpha _B} in the above equation.

θB(t)=θ0+(ω02)(tt1)+α(tt1)2{\theta _{\rm{B}}}\left( t \right) = {\theta _0} + \left( {\frac{{{\omega _0}}}{2}} \right)\left( {t - {t_1}} \right) + \alpha {\left( {t - {t_1}} \right)^2}

(H)

The angular position of particle A is,

θA(t)=θ0+ω0t+12αt2{\theta _{\rm{A}}}\left( t \right) = {\theta _0} + {\omega _0}t + \frac{1}{2}\alpha {t^2}

Differentiate the above equation with respect to tt.

ωA(t)=ω0+αt{\omega _{\rm{A}}}\left( t \right) = {\omega _0} + \alpha t …… (1)

The angular position of particle B is,

θB(t)=θ0+(ω02)(tt1)+α(tt1)2{\theta _{\rm{B}}}\left( t \right) = {\theta _0} + \left( {\frac{{{\omega _0}}}{2}} \right)\left( {t - {t_1}} \right) + \alpha {\left( {t - {t_1}} \right)^2}

Differentiate the above equation with respect to tt.

ωB(t)=(ω02)+2α(tt1){\omega _{\rm{B}}}\left( t \right) = \left( {\frac{{{\omega _0}}}{2}} \right) + 2\alpha \left( {t - {t_1}} \right) …… (2)

Equate equations (1) and (2).

(ω02)+2α(tt1)=ω0+αt\left( {\frac{{{\omega _0}}}{2}} \right) + 2\alpha \left( {t - {t_1}} \right) = {\omega _0} + \alpha t

Solve the equation(ω02)+2α(tt1)=ω0+αt\left( {\frac{{{\omega _0}}}{2}} \right) + 2\alpha \left( {t - {t_1}} \right) = {\omega _0} + \alpha tfor t.

t=ω0+2αt12αt = \frac{{{\omega _0} + 2\alpha {t_1}}}{{2\alpha }}

Ans: Part A

The quantity represented by θ\theta is a function of time.

Part B

The quantity represented by θ0{\theta _0} is not a function of time.

Part C

The quantity represented by ω0{\omega _0} is not a function of time.

Part D

The quantity represented by ω\omega is a function of time.

Part E

The equation ω2=ω02+2α(θθ0){\omega ^2} = {\omega _0}^2 + 2\alpha \left( {\theta - {\theta _0}} \right)is not an explicit function of time.

Part F

The time variable tt represents the time elapsed from when the angular velocity equalsω0{\omega _0} until the angular velocity equals ω\omega .

Part G

The equation, θB(t)=θ0+(ω02)(tt1)+α(tt1)2{\theta _{\rm{B}}}\left( t \right) = {\theta _0} + \left( {\frac{{{\omega _0}}}{2}} \right)\left( {t - {t_1}} \right) + \alpha {\left( {t - {t_1}} \right)^2}describes the angular position of particle B.

Part H

The time after which the angular velocity of particle A will be equal to the angular velocity of particle B is, t=ω0+2αt12αt = \frac{{{\omega _0} + 2\alpha {t_1}}}{{2\alpha }}.